OpenStudy (anonymous):

A tennis ball is thrown straight up into the air with an initial velocity of 22.5 m/s. It is caught at the same distance above the ground. How high does the ball rise?

OpenStudy (jfraser):

use the equations of motion, and solve for a vertical displacement\[V_f = V_i + a*t\]or\[V_f^2 = V_i^2 + 2*a*\Delta y\]or\[\Delta y = V_i*t + \frac{1}{2}*a*t^2\]